# Rotate, Clip, and Partition: Towards W2A4KV4 Quantization by Integrating Rotation and Learnable Non-uniform Quantizer

### v1

@@ 5.3 Ablation Studies
Table 5: Ablation study on the impact of each component of our proposed RCP on performance for LLaMA-2 7B.
@@
As shown in Table 5, we conducted an ablation study to analyze the impact of removing each component of RCP on model performance. In 4-bit activation quantization, addressing the outliers in activations was crucial, and this was effectively resolved using rotation, which led to the largest performance gain compared to the baseline. This demonstrates that rotation is a viable solution when quantizing activations to low bit-width.
However, we found that the narrow weight distribution caused by rotation hindered successful training of LWC. Specifically, when examining the training process with rotation applied during LWC training, the training loss curve exhibited instability. The combination of low bit-width quantization challenges and the difficulty in finding an optimal LWC required training stabilization, which was achieved by LDP. LDP reduced PPL from 10.59 to 8.31, demonstrating a clear advantage.
@@
Since the rotation requires additional processes before and after inference, we investigated the performance trend by incrementally adding rotation matrices ($R_{1}$,$R_{2}$,$R_{3}$,$R_{4}$) to different components to find an appropriate balance between accuracy and overhead. The results are presented in Table 6. The table demonstrates that the impact of the rotation was most significant with $R_{1}$ and $R_{4}$. Especially, $R_{1}$, which applies rotation matrix to the input weight and input activation of all modules thereby having the largest impact on quantization performance. Additionally, our analysis revealed that in LLaMA-2 7B, the input to the down projection layer (of the MLP) exhibited a significant number of outliers, which was effectively addressed through $R_{4}$ online rotation to activation.
@@ A.1 Related Works
GPTQ Frantar et al. (2022) introduced an accurate post-training quantization (PTQ) method based on approximate second-order information that enables weight-only quantization down to 3-4 bits through block-wise reconstruction. SmoothQuant Xiao et al. (2023) proposed smoothing activation outliers by offline migrating quantization difficulty from activations to weights through equivalent transformation, enabling accurate 8-bit weight-activation quantization. AWQ Lin et al. (2024a) built upon SmoothQuant’s equivalent transformation concept but introduced activation-aware channel-wise scaling to protect salient weights during weight-only quantization. OmniQuant Shao et al. (2024) enhanced quantization by introducing learnable weight clipping and equivalent transformation parameters that are jointly optimized through block-wise reconstruction.
LLM-QAT Liu et al. (2024a) was the first to explore quantization-aware training (QAT) for LLMs using data-free knowledge distillation from the full-precision model to guide low-bit quantization. BitDistiller Du et al. (2024) improved upon LLM-QAT by introducing a self-distillation framework with confidence-aware KL divergence to enable sub-4-bit quantization while maintaining efficiency. EfficientQAT Chen et al. (2024b) made QAT more practical by introducing block-wise training of all parameters followed by end-to-end training of quantization parameters.
@@
QuaRot Ashkboos et al. (2024b) introduced a rotation-based approach using Hadamard transformations to eliminate outliers in activations and KV-cache, enabling end-to-end 4-bit quantization including weights, activations and KV-cache. SpinQuant Liu et al. (2024b) enhanced this rotation-based approach by learning optimal rotation matrices instead of using random ones.
@@
Atom Zhao et al. (2024) first introduced W4A4 quantization for LLM serving but faced performance challenges from dequantization overhead. QServe Lin et al. (2024b) addressed the challenges by introducing W4A8KV4 quantization with progressive group quantization FLUTE Guo et al. (2024) focused on developing efficient GPU kernels for flexible lookup table-based quantization methods that can support arbitrary bit widths including 3-bit and 4-bit quantization.
A.2 Reasoning Task Example: HumanEval
A.3 Implementation Details
@@
A.4 More Ablation Studies
@@
We investigate the impact of preserving the orthogonality of the rotations by modifying the LLaMA-2 model implementation to apply all rotation operators online while freezing the rotation matrices. Table 7 presents the results. Applying factorized rotation prevents the fusion of the rotation matrix into the weight tensor, resulting in an increase in the number of intermediate tensors (rotation matrix and intermediate activation), which significantly raises VRAM requirements. For instance, applying only $R_{1}$ needs to reduce the training batch size from 8 to 1. Under the condition of maintaining an equal total number of tokens processed by the model, we compared the performance of W2A16KV16 with only $R_{1}$ applied. The perplexity of BitDistiller with $R_{1}$ fused was 7.6, whereas applying QAT with factorized rotation resulted in a PPL of 12.5. This indicates that performing weight updates through QAT while preserving $R1$ orthogonality hinders QAT optimization. This is because the factorization constrains the weight updates to a restricted space defined by the factorized condition, requiring the backpropagation process to maintain within this space. This limitation reduces the flexibility of optimization, making it challenging to efficiently adjust the weights. Consequently, this leads to suboptimal training dynamics and ultimately results in degraded model performance. Furthermore, extending factorization to $R_{2}$ and $R_{4}$ would lead to an even greater increase in VRAM usage. In contrast, training fused weight effectively alters only the distribution and is analogous to standard LLM training, which is well-known to perform effectively. In summary, given that resource consumption increases while performance degrades, we have decided not to explicitly preserve orthogonality and instead allow the algorithm to handle this aspect.
@@
Recent work introduced layerwise QAT Chen et al. (2024b), which updates one layer at a time while freezing others, allowing training on a single GPU. We extended this approach by applying rotation but observed significant performance degradation. The main issue stemmed from fuse rotation matrices in the weights; layerwise updates disrupted orthogonality, preventing the activation space from restoring its original space, leading to cumulative errors and reduced accuracy. In contrast, end-to-end methods like BitDistiller naturally mitigate this issue during updates. While factorized rotation could help, its high GPU memory requirements for holding rotation matrices and intermediate tensors on GPU memory offsets the advantage. Despite these challenges, exploring single GPU training using rotation matrix remains a promising direction for future work.
A.5 GEMM Kernel Design for Non-uniform W2A4 Quantization

### v2

@@ 5.3 Ablation Studies
Table 9: Ablation study on the impact of each component of RCP on performance for LLaMA-2 7B.
@@
As shown in Table 9, we conducted an ablation study to analyze the impact of removing each component of RCP on model performance. In 4-bit activation quantization, addressing the outliers in activations was crucial, and this was effectively resolved using rotation, which led to the largest performance gain compared to the baseline. This demonstrates that rotation is a viable solution when quantizing activations to low bit-width. Learning the clipping range with LWC (Rotation $\checkmark$ and LWC $\checkmark$) gives a small extra reduction in PPL. This shows that the initial range computed by the grid search algorithm of AWQ is already reasonable, but adjusting the clipping parameters to better fit the updated weight can be beneficial for QAT. On top of these two components, LDP further reduces quantization error under extremely low bit-width settings.
@@
Since the rotation requires additional processes before and after inference, we investigated the performance trend by incrementally adding rotation matrices ($R_{1}$$R_{2}$$R_{3}$$R_{4}$) to different components to find an appropriate balance between accuracy and overhead. The results are presented in Table 10. The table demonstrates that the impact of the rotation was most significant with $R_{1}$ and $R_{4}$. Especially, $R_{1}$, which applies rotation matrix to the input weight and input activation of all modules thereby having the largest impact on quantization performance. Additionally, our analysis revealed that in LLaMA-2 7B, the input to the down projection layer (of the MLP) exhibited a significant number of outliers, which was effectively addressed through $R_{4}$ online rotation to activation. Additional ablation results can be found in Appendix A.5.
@@ 6 Conclusion
Acknowledgments
This work was supported by Samsung Advanced Institute of Technology, and MX division, Samsung Electronics Co., Ltd., Inter-university Semiconductor Research Center (ISRC) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO. RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)].
@@ A.1 Related Works
GPTQ Frantar et al. (2022) introduced an accurate post-training quantization (PTQ) method based on approximate second-order information that enables weight-only quantization down to 3-4 bits through block-wise reconstruction. SmoothQuant Xiao et al. (2023) proposed smoothing activation outliers by offline migrating quantization difficulty from activations to weights through equivalent transformation, enabling accurate 8-bit weight-activation quantization. AWQ Lin et al. (2024b) built upon SmoothQuant’s equivalent transformation concept but introduced activation-aware channel-wise scaling to protect salient weights during weight-only quantization. OmniQuant Shao et al. (2024) enhanced quantization by introducing learnable weight clipping and equivalent transformation parameters that are jointly optimized through block-wise reconstruction.
LLM-QAT Liu et al. (2024a) was the first to explore quantization-aware training (QAT) for LLMs using data-free knowledge distillation from the full-precision model to guide low-bit quantization. BitDistiller Du et al. (2024) improved upon LLM-QAT by introducing a self-distillation framework with confidence-aware KL divergence to enable sub-4-bit quantization while maintaining efficiency. EfficientQAT Chen et al. (2024) made QAT more practical by introducing block-wise training of all parameters followed by end-to-end training of quantization parameters.
@@
QuaRot Ashkboos et al. (2024b) introduced a rotation-based approach using Hadamard transforms to eliminate outliers in activations and KV-cache, enabling end-to-end 4-bit quantization including weights, activations and KV-cache. SpinQuant Liu et al. (2024b) enhanced this rotation-based approach by learning optimal rotation matrices instead of using random ones.
@@
Atom Zhao et al. (2024) first introduced W4A4 quantization for LLM serving but faced performance challenges from dequantization overhead. QServe Lin et al. (2024c) addressed the challenges by introducing W4A8KV4 quantization with progressive group quantization FLUTE Guo et al. (2024) focused on developing efficient GPU kernels for flexible lookup table-based quantization methods that can support arbitrary bit widths including 3-bit and 4-bit quantization.
A.2 Proof of Lemma 1
Proof of Lemma 1.
The proof follows directly from Sub-lemmas 1.11.4. Specifically, Sublemma 1.4 shows that for all components $Y_{i}$ of the transformed vector:
$\text{Kurt}(Y_{i})=\frac{\text{Kurt}(X)}{n}$
Since $\text{Kurt}(X)<0$ for platykurtic distributions and $n>1$:
$\frac{\text{Kurt}(X)}{n}>\text{Kurt}(X)$
Therefore:
$\text{Kurt}(Y_{i})>\text{Kurt}(X)$
for all $i\in\{1,2,\ldots,n\}$, which completes the proof. ∎
Definition.
The $n\times n$ normalized Hadamard matrix $\mathbf{H}_{n}$, where $n=2^{k}$ for some non-negative integer $k$, is defined as:
$\mathbf{H}_{n}=\frac{1}{\sqrt{n}}\mathbf{H}^{\prime}_{n}$
where $\mathbf{H}^{\prime}_{n}$ is the unnormalized Hadamard matrix with elements $H^{\prime}_{ij}\in\{-1,1\}$ constructed recursively as:
$\mathbf{H}^{\prime}_{1}=[1],\quad\mathbf{H}^{\prime}_{2^{k+1}}=\begin{bmatrix}\mathbf{H}^{\prime}_{2^{k}}&\mathbf{H}^{\prime}_{2^{k}}\\ \mathbf{H}^{\prime}_{2^{k}}&-\mathbf{H}^{\prime}_{2^{k}}\end{bmatrix}$
Note that the first row of $\mathbf{H}^{\prime}_{n}$ consists entirely of 1s, while every other row contains exactly $n/2$ entries of 1 and $n/2$ entries of 1-1.
Furthermore, the normalized Hadamard matrix $\mathbf{H}_{n}$ is orthogonal:
$\mathbf{H}_{n}\mathbf{H}_{n}^{T}=\mathbf{I}_{n}$
where $\mathbf{I}_{n}$ is the $n\times n$ identity matrix. For any random vector $\mathbf{X}$ with independent components of identical variance $\sigma^{2}$, the transformed vector $\mathbf{Y}=\mathbf{H}_{n}\mathbf{X}$ has the same component-wise variance:
$\text{Var}(Y_{i})=\sigma^{2}\quad\text{for all }i\in\{1,2,\ldots,n\}$ (15)
This follows from the fact that for a covariance matrix $\Sigma_{X}=\sigma^{2}\mathbf{I}_{n}$, the transformed covariance is $\Sigma_{Y}=\mathbf{H}_{n}\Sigma_{X}\mathbf{H}_{n}^{T}=\sigma^{2}\mathbf{H}_{n}\mathbf{H}_{n}^{T}=\sigma^{2}\mathbf{I}_{n}=\Sigma_{X}$.
Sublemma 1.1.
Each component $Y_{i}$ of the transformed vector can be expressed as a linear combination of the original variables:
$Y_{i}=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}H^{\prime}_{ij}X_{j}$ (16)
where $H^{\prime}_{ij}\in\{-1,1\}$ are the elements of the unnormalized Hadamard matrix.
Proof.
By definition of matrix multiplication, each component $Y_{i}$ of the transformed vector $\mathbf{Y}=\mathbf{H}_{n}\mathbf{X}$ is given by:
$Y_{i}=\sum_{j=1}^{n}h_{ij}X_{j}=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}H^{\prime}_{ij}X_{j}$
where $h_{ij}=\frac{H^{\prime}_{ij}}{\sqrt{n}}$ are the elements of the normalized Hadamard matrix. ∎
Sublemma 1.2.
If $X_{i}$ has mean $\mu$ and variance $\sigma^{2}$, then $Y_{i}$ has mean $\mu_{Y}$ and variance $\sigma_{Y}^{2}$ where:
$\mu_{Y}=\begin{cases}\sqrt{n}\mu&\text{if }i=1\\ 0&\text{if }i\neq 1\end{cases}$
$\sigma_{Y}^{2}=\sigma^{2}$
Proof.
Let’s calculate the mean of each transformed component $Y_{i}$:
$\begin{split}\mathbb{E}[Y_{i}]=\mathbb{E}\left[\frac{1}{\sqrt{n}}\sum_{j=1}^{n}H^{\prime}_{ij}X_{j}\right]\\ =\frac{1}{\sqrt{n}}\sum_{j=1}^{n}H^{\prime}_{ij}\mathbb{E}[X_{j}]=\frac{\mu}{\sqrt{n}}\sum_{j=1}^{n}H^{\prime}_{ij}\end{split}$
For $i=1$, the first row of the unnormalized Hadamard matrix consists entirely of 1s. Therefore:
$\mathbb{E}[Y_{1}]=\frac{\mu}{\sqrt{n}}\cdot n=\sqrt{n}\mu$
For all other rows $i>1$, the Hadamard matrix has the property that each row contains exactly $\frac{n}{2}$ entries of 1 and $\frac{n}{2}$ entries of -1. This gives:
$\mathbb{E}[Y_{i}]=\frac{\mu}{\sqrt{n}}\left(\frac{n}{2}-\frac{n}{2}\right)=0\quad\text{for }i>1$
For the variance, assuming independence of $X_{j}$:
$\begin{split}\text{Var}(Y_{i})=\text{Var}\left(\frac{1}{\sqrt{n}}\sum_{j=1}^{n}H^{\prime}_{ij}X_{j}\right)\\ =\frac{1}{n}\sum_{j=1}^{n}(H^{\prime}_{ij})^{2}\cdot\text{Var}(X_{j})\end{split}$
Since $(H^{\prime}_{ij})^{2}=1$ for all $i,j$ and all $X_{j}$ have variance $\sigma^{2}$:
$\text{Var}(Y_{i})=\frac{\sigma^{2}}{n}\cdot n=\sigma^{2}\quad\text{for all }i$
Sublemma 1.3 (Relationship Between Fourth Moments).
For a sum of independent random variables with identical distributions, the standardized fourth cumulant (excess kurtosis) of the sum relates to the individual excess kurtosis by:
$\text{Kurt}\left(\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\epsilon_{j}X_{j}\right)=\frac{\text{Kurt}(X)}{n}$
where $\epsilon_{j}\in\{-1,1\}$ and $X_{j}$ are i.i.d. with the same distribution as $X$.
Proof.
Let $Z=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\epsilon_{j}X_{j}$ where $\epsilon_{j}\in\{-1,1\}$ and $X_{j}$ are i.i.d. with the same distribution as $X$.
The excess kurtosis of a random variable $W$ is defined as:
$\text{Kurt}(W)=\frac{\mathbb{E}[(W-\mathbb{E}[W])^{4}]}{(\text{Var}(W))^{2}}-3$
For independent random variables, the cumulants of a sum equal the sum of the cumulants. The fourth cumulant $\kappa_{4}$ corresponds to:
$\begin{split}\kappa_{4}(W)=\mathbb{E}[(W-\mathbb{E}[W])^{4}]-3(\mathbb{E}[(W-\mathbb{E}[W])^{2}])^{2}\\ =\text{Var}(W)^{2}\cdot\text{Kurt}(W)\end{split}$
For our sum $Z$, the fourth cumulant is:
$\kappa_{4}(Z)=\sum_{j=1}^{n}\kappa_{4}\left(\frac{\epsilon_{j}X_{j}}{\sqrt{n}}\right)$
Since $\kappa_{4}(\alpha X)=\alpha^{4}\kappa_{4}(X)$ for any scalar $\alpha$:
$\kappa_{4}(Z)=\sum_{j=1}^{n}\frac{\epsilon_{j}^{4}}{n^{2}}\kappa_{4}(X_{j})=\frac{1}{n^{2}}\sum_{j=1}^{n}\kappa_{4}(X_{j})$
Given that $\epsilon_{j}^{4}=1$ and all $X_{j}$ have the same distribution:
$\kappa_{4}(Z)=\frac{n}{n^{2}}\kappa_{4}(X)=\frac{\kappa_{4}(X)}{n}$
Since $\kappa_{4}(X)=\text{Var}(X)^{2}\cdot\text{Kurt}(X)$ and $\text{Var}(Z)=\text{Var}(X)$ (as shown in Sublemma 1.2):
$\text{Var}(Z)^{2}\cdot\text{Kurt}(Z)=\frac{\text{Var}(X)^{2}\cdot\text{Kurt}(X)}{n}$
Therefore:
$\text{Kurt}(Z)=\frac{\text{Kurt}(X)}{n}$
Sublemma 1.4 (Application to Hadamard Transform).
For a random vector with i.i.d. components having negative excess kurtosis ($\text{Kurt}(X)<0$), after applying the normalized Hadamard transform:
$\text{Kurt}(Y_{i})=\frac{\text{Kurt}(X)}{n}$
Since $\text{Kurt}(X)<0$ and $n>1$, we have $\frac{\text{Kurt}(X)}{n}>\text{Kurt}(X)$, which implies $\text{Kurt}(Y_{i})>\text{Kurt}(X)$.
Proof.
From Sublemma 1.1, each component $Y_{i}$ of the Hadamard transform can be written as:
$Y_{i}=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}H^{\prime}_{ij}X_{j}$
This is precisely the form analyzed in Sublemma 1.3, with $\epsilon_{j}=H^{\prime}_{ij}$.
For $i>1$ (where the mean is 0), applying Sublemma 1.3 directly:
$\text{Kurt}(Y_{i})=\frac{\text{Kurt}(X)}{n}$
For $i=1$, we need to account for the non-zero mean. We can center the variable:
$Y_{1}-\mathbb{E}[Y_{1}]=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}(X_{j}-\mu)$
Applying the same cumulant analysis to this centered variable:
$\text{Kurt}(Y_{1})=\frac{\text{Kurt}(X)}{n}$
Since we assumed $\text{Kurt}(X)<0$ for a platykurtic distribution, and $n>1$:
$\frac{\text{Kurt}(X)}{n}>\text{Kurt}(X)$
Therefore, for all components $i\in\{1,2,\ldots,n\}$:
$\text{Kurt}(Y_{i})>\text{Kurt}(X)$
A.3 Additional Results on the Kurtosis Analysis
Details on Fig. 3
Let $T$, $C$, and $H$ denote the sequence length of an input, the input channel of weight, and the output channel of weight, respectively. Since RCP works on group-wise quantization, we additionally denote the group size as $G$ and the number of groups as $N$ so that $C=NG$. The input activation $\mathbf{X}$ and its rotated version $\mathbf{X}_{r}$ both have a dimension of $(T,C)$, and the weight $\mathbf{W}$ and its rotated & clipped version $\mathbf{W}_{rc}$ a dimension of $(C,H)$.
We compute the group-wise excess kurtosis $Kurt_{group}$ by reshaping the weights into $(N,G,H)$, and computing the excess kurtosis along the second dimension, resulting in a shape $(N,H)$. For the mean absolute error $QErr(\mathbf{W},\mathbf{X})$, the shape of the output activation $\mathbf{XW}$ and the quantized version $\mathbf{X}Q(\mathbf{W})$ is $(T,H)$. Then, we average $Kurt_{group}$ along the $N$ dimension so that the excess kurtosis values of the quantization groups contributing to a single output activation element are averaged. The mean absolute error of the output activation is averaged along the $T$ dimension to measure the mean increase of the quantization error in each output activation element.
In all two-dimensional histogram plots, the range is limited to $1.5\sigma,1.5\sigma]$ for both axes to prevent outliers from occupying most of the space.
More Plots and Discussion
We repeat the same experiments as in Section 3 on three different transformer layers (0, 15, and 31) and three different types of weights (q_projq\_proj , o_projo\_proj, and down_projdown\_proj) and the plots are presented in Fig. 8. In each subplot’s title, we specify the ratio of quantization groups that are platykurtic (i.e., $p_{platy}=Kurt(\mathbf{W}_{group})<0$. When $p_{platy}\ll 0.5$ (Fig. 8(d) and 8(g)), the Hadamard transform decreases the excess kurtosis, possibly reducing the average quantization error 8(d). When $p_{platy}\approx 0.5$, the excess kurtosis is increased in almost all cases, with the average quantization error also enlarged. This supports our claim in Lemma 1 and Section 3 empirically.
W2A4KV4 Loss Curve Comparison
Figure 6: Comparison of training loss curves for LLaMA-2 7B under W2A4KV4 quantization: BitDistiller + Rotate, BitDistiller + Rotate + LWC, and RCP.
Figure 6 presents the training loss curves of RCP and BitDistiller with rotation, and BitDistiller with rotation and LWC during W2A4KV4 quantization. While all methods eventually converge, a notable performance gap originates from the early stages of training. As highlighted in the magnified region ($\leq 100$ steps), BitDistiller with rotation exhibits pronounced loss spikes, indicating unstable optimization. Even with the addition of LWC, such loss spikes persist, ultimately leading to suboptimal final performance.
In contrast, RCP, which incorporates Learnable Direct Partitioning (LDP), demonstrates significantly more stable loss curve from the beginning of training. These results provide evidence for the necessity of LDP when applying extreme low-bit quantization with rotation and clipping.
A.4 Additional Experimental Results
Implementation Details
Tensor Type Sym / Asym Group Quant. Clipping Ratio
Weight Asym Yes (Size=128) LWC
Activation Sym No (Per-token) 0.9
KV-cache Asym Yes (Size=128) 0.95
Table 11: Quantization configurations for each component.
@@ A.4 Additional Experimental Results
As shown in Table 11, we apply an asymmetric LDP with LWC to weights, a symmetric uniform quantizer to activations, and an asymmetric uniform quantizer with a group size of 128 to the KV-cache, with clipping ratios of 0.9 and 0.95 for activations and KV-cache, respectively.
Additional GEMV Benchmarks
To compare the gain solely attributed to our non-uniform W2A4 GEMV kernel, we also apply the inefficient quantizer and the online transform to FP16 weights so that the W16A4 model is simulated, and the measured latency values are listed in Table 17. Using online FP16 Hadamard transform, our RCP GEMV is faster than PyTorch nn.Linear kernel, which indicates that our GEMV implementation is faster and can successfully hide its latency to the following activation quantization.
A.5 Additional Ablation Studies
@@
We investigate the impact of preserving the orthogonality of the rotations by modifying the LLaMA-2 model implementation to apply all rotation operators online while freezing the rotation matrices. Table 12 presents the results. Applying factorized rotation prevents the fusion of the rotation matrix into the weight tensor, resulting in an increase in the number of intermediate tensors (rotation matrix and intermediate activation), which significantly raises VRAM requirements. For instance, applying only $R_{1}$ needs to reduce the training batch size from 8 to 1. Under the condition of maintaining an equal total number of tokens processed by the model, we compared the performance of W2A16KV16 with only $R_{1}$ applied. The perplexity of BitDistiller with $R_{1}$ fused was 7.6, whereas applying QAT with factorized rotation resulted in a PPL of 12.5. This indicates that performing weight updates through QAT while preserving $R_{1}$ orthogonality hinders QAT optimization. This is because the factorization constrains the weight updates to a restricted space defined by the factorized condition, requiring the backpropagation process to maintain within this space. This limitation reduces the flexibility of optimization, making it challenging to efficiently adjust the weights. Consequently, this leads to suboptimal training dynamics and ultimately results in degraded model performance. Furthermore, extending factorization to $R_{2}$ and $R_{4}$ would lead to an even greater increase in VRAM usage. In contrast, training fused weight effectively alters only the distribution and is analogous to standard LLM training, which is well-known to perform effectively. In summary, given that resource consumption increases while performance degrades, we have decided not to explicitly preserve orthogonality and instead allow the algorithm to handle this aspect.
@@
Recent work introduced layerwise QAT Chen et al. (2024), which updates one layer at a time while freezing others, allowing training on a single GPU. We extended this approach by applying rotation but observed significant performance degradation. The main issue stemmed from fusing rotation matrices in the weights; layerwise updates disrupted orthogonality, preventing the activation space from restoring its original space, leading to cumulative errors and reduced accuracy. In contrast, end-to-end methods like BitDistiller naturally mitigate this issue during updates. While factorized rotation could help, its high GPU memory requirements for holding rotation matrices and intermediate tensors on GPU memory offsets the advantage. Despite these challenges, exploring single GPU training using rotation matrix remains a promising direction for future work.
A.6 GEMM Kernel Design for Non-uniform W2A4 Quantization